in a triangle ABC, angle B is 5 times angle A and angle C is 9 degrees less than 4 times angle A.
Find the size of the angles ... please

Answer by Earlsdon(6294)

Start with A+B+C = 180, then:
B = 5A "Angle B is 5 times angle A".
C = 4A-9 "Angle C is 9 degrees less than 4 times angle A".
Now substitute these values into the first equation to get:
A+5A+4A-9 = 180 Now simplify and solve for A.
10A-9 = 180 Add 9 to both sides.
10A = 189 Divide both sides by 10.
A = 18.9 degrees.
B = 5A
B = 5(18.9)
B = 94.5 degrees.
C = 4A-9
C = 4(18.9)-9
C = 75.6-9
C = 66.6 degrees.
Check:
A+B+C = 180
18.9+94.5+66.6 = 180 OK